OpenStudy (anonymous):

Find the distance between points P(8, 2) and Q(3, 8) to the nearest tenth.

OpenStudy (wolf1728):

x distance = 8-3 = 5 y distance = 2-8 = 6 By Pythagorean Theorem hypotenuse² = x² + y² hypotenuse² = 25 + 36 hypotenuse = sq root (61) hypotenuse = 7.81 hypotenuse = distance between the 2 points

OpenStudy (anonymous):

Yeah, thanks for the answer. I figured that I had to use the Distance formula.

OpenStudy (wolf1728):

Here's a graph to double check the calculation.

OpenStudy (wolf1728):

Okay you are welcome :-)